OpenStudy (anonymous):

A triangular playground has angles with measures in the ratio 8 : 4 : 3. What is the measure of the smallest angle?

OpenStudy (anonymous):

let the angles be 8x, 4x, 3x \[\huge 8x+ 4x+ 3x=180^0\] \[\huge 15x=180^0 \rightarrow x = \frac{180^0}{15} = 12^0\] Smallest angle = 3x = 3 * 12 = 36 degrees

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

Can you help me with a few more? I was absent.

OpenStudy (anonymous):

Sure, go ahead..

OpenStudy (anonymous):

Write an in slope-intercept form of the line through point P(-8,6) with slope 4.

OpenStudy (anonymous):

Ok, pls wait..

OpenStudy (anonymous):

Given: \[\huge (x_1, y_1) = (-8,6) \rightarrow x_1=-8, y_1= 6\] Slope m= 4The eq. of line in slope-point form is given as: \[\huge (y- y_1) = m((x-x_1)\] Substituting the above values in the eq we find: \[\huge (y- 6) = 4[x-(-8)] \rightarrow (y- 6) = 4(x+8)\] \[\huge \rightarrow (y- 6) = 4(x+8) \rightarrow y-6= 4x- 32\] \[\huge \rightarrow y= 4x- 32+6\rightarrow y= 4x- 26\]

OpenStudy (anonymous):

@BlveXO

OpenStudy (anonymous):

Okay, just one more. :P

OpenStudy (anonymous):

You are a tremendous help.

OpenStudy (anonymous):

What is an equation for the line that passes through points (2,-7) and (4,3)?

OpenStudy (anonymous):

Since line is passing through points (2,-7) and (4,3), so eq of line in 2-point form is given as: Here: (x1, y1) =(2,-7) & (x2, y2) = (4,3) \[\huge {y-y_1} = m(x-x_1)\] \[ \huge \rightarrow {y-y_1} = \frac{(y_2-y_1)}{(x_2-x_1)}(x-x_1)\] Now susstituting the above values we find: \[ \huge \rightarrow {y-(-7)} = \frac{3-(-7)}{4-2}(x-2)\] \[ \huge \rightarrow {y+7} = \frac{3+7}{2}(x-2)\] \[ \huge \rightarrow {y+7} = \frac{10}{2}(x-2)\] \[ \huge \rightarrow {y+7} = 5(x-2)\] \[ \huge \rightarrow {y+7} = 5x-10\] \[ \huge \rightarrow {y} = 5x-10-7\] \[ \huge \rightarrow y= 5x-17\] is the required eq of the line.

OpenStudy (anonymous):

@BlveXO

OpenStudy (anonymous):

Okay, thank you.! :)